On April 1, Sangvikar Company had the following balances in its inventory accounts:

Materials Inventory $12,990
Work-in-Process Inventory 21,220
Finished Goods Inventory 8,840

Work-in-process inventory is made up of three jobs with the following costs:

Job 114 Job 115 Job 116
Direct materials $2,403 $2,692 $3,103
Direct labor   1,780   1,560   4,320
Applied overhead   1,246   1,092 3.024

During April, Sangvikar experienced the transactions listed below.

  1. Materials purchased on account, $29,050.

  2. Materials requisitioned: Job 114, $16,060; Job 115, $12,220; and Job 116, $5,470.

  3. Job tickets were collected and summarized: Job 114, 140 hours at $12 per hour; Job 115, 210 hours at $15 per hour; and Job 116, 100 hours at $17 per hour.

  4. Overhead is applied on the basis of direct labor cost.

  5. Actual overhead was $4,655.

  6. Job 115 was completed and transferred to the finished goods warehouse.

  7. Job 115 was shipped, and the customer was billed for 125 percent of the cost.

Instructions:

1) Calculate the predetermined overhead rate based on direct labor cost

2) Calculate the ending balance for each job as of April 30

3) Calculate the ending balance of work-in-process as of April 30

4) Calculate the cost of Goods Sold for April

5) Assuming that Sangvikar prices its jobs at cost plus 25 percent, calculate the price of the one job that was sold during April (round to the nearest dollar)
Answer #1

1)Beginning applied overhead to beginning work in process inventory = 1246+1092+3024=5362

Direct labor cost ,Beginning= 1780+1560+4320=7660

Predetermined overhead rate = Applied overhead /Direct labor cost

                                = 5362 /7660

                                = .70 or 70%

2)

114 115 116
Beginning work in process inventory 5429 5344 10447
current period cost
Direct material 16060 12220 5470
Direct labor 140*12=1680 210*15=3150 100*17=1700
overhead applied 1680*70%= 1176 3150*70%= 2205 1700*70%= 1190
Total cost 24345 22919 18807

3)Ending work in process = cost of Job 114 + 116

                          = 24345 +18807

                         = $ 43152

4)cost of goods sold = Job 115 = 22919

5)Price = cost (1+ % on cost)

              = 22919 (1+.25)

              = $ 28648.75 (rounded to 28649)